OpenStudy (anonymous):

Suppose that 3<=f'(x)<=5 for all values of x. Show that 18<=f(8)-f(2)<=30

OpenStudy (anonymous):

\[3\le f'(x)\le5\]

OpenStudy (anonymous):

\[18\le f(8)-f(2)\le30\]

OpenStudy (zarkon):

mean value theorem

OpenStudy (anonymous):

which is

OpenStudy (zarkon):

look it up...you need it to do this problem

OpenStudy (anonymous):

i have the theorem bu ti don't get how it applies

OpenStudy (zarkon):

look at \[\frac{f(8)-f(2)}{8-2}=f'(c)\] \[2<c<8\]

OpenStudy (anonymous):

i still don't get how that helps me find 18≤f(8)−f(2)≤30

OpenStudy (zarkon):

think about it for a little bit...if I do any more I will be doing the entire problem

OpenStudy (anonymous):

i am not good at math. i never understand it